# Spark Transformer: Reactivating Sparsity in FFN and Attention

### v1

@@ 4.1 Quality
We evaluate Spark Transformer on a suite of benchmarks that are used in the Gemma-2 paper [27], and report the result in Figure 1(b). We observe that Spark Transformer matches the quality of Gemma-2 while having a drastically reduced FLOP count per token.
The near-quality neutrality of Spark Transformer distinguishes it from related work on enforcing activation sparsity in FFN, which often lead to a quality loss. To demonstrate this, we pretrain variants of Gemma-2 2B with 1) activation function switched to ReLU [86], 2) activation function switched to ReLU2 [59], and 3) Top-k thresholding applied before GELU. Due to high training cost, all models are pretrained for 1/6 of the standard Gemma-2 training iterations. In Figure 1(a), we report the training losses and FLOPs per token relative to those of the standard Gemma-2. It can be seen that these models either suffer from a large quality loss (i.e., ReLU and Topk), or do not lead to a sufficient FLOPs reduction (i.e., ReLU2). In contrast, our Spark FFN achieves less quality loss with more FLOPs reduction. [^3] Notably, the only difference between Spark FFN and Topk, when both have 8% activated neurons, is in the introduction of the low-cost predictor. Hence, this predictor not only leads to FLOPs reduction but also (surprisingly?) improves model quality. Finally, the combination of Spark FFN with Spark Attention introduces additional FLOPs reduction and quality benefits, enabling an almost neutral quality of Spark Transformer with a large overall FLOPs reduction.
@@ 4.2 Sparsity
To verify the effectiveness of statistical top-$k$, we report the level of sparsity measured in terms of percentage of nonzeros in FFN and the number of nonzeros in Attention. At the beginning of model training, we observe that statistical top-$k$ produces close to 8% nonzeros in FFN (see Figure 3(a)), which aligns well with our hyper-parameter choice of using $k/d_{\text{ff}}=8\%$ in Spark FFN. This is expected as the model parameters, particularly ${\bm{K}}$ in Spark FFN, are randomly initialized, hence the entries of the activation maps are drawn from a Gaussian distribution which is in accordance with the assumption of statistical top-$k$. The Gaussian assumption is no longer guaranteed after training, but we empirically observe it to hold approximately (see Section C.1) and statistical top-$k$ reliably produce a sparsity level close to 8% until the end of training at $480k$ steps. Sparsity in attention is reported in Figure 3(b), which show that the number of attended tokens is below our hyper-parameter choice of 256 in Spark Attention throughout training. In particular, the numbers are much smaller because the results are from averaging over all tokens many of which have a context length of less than 256. Finally, we observe comparable levels of sparsity during evaluation (see Section C.2).
@@ 4.3 Inference Efficiency
We evaluate the efficiency benefits of Spark Transformer over standard Gemma-2 on both CPUs and GPUs. For CPU evaluation, we use gemma.cpp [30], the official C++ inference engine optimized for CPUs. For GPU evaluation, we use llama.cpp [28], a widely-used LLM inference engine which supports running a wide selection of LLM models on GPUs. We modify both frameworks to support sparse matrix multiplication operators, which exploit sparsity in both FFN and Attention layers. Our implementation leverages vector SIMD operations [75] existing in modern CPU, and customized CUDA kernel for GPU. See Figure 4 for an illustration and Section B in Appendix for details. We show that Spark Transformer significantly improves the efficiency of transformer models, even in highly FLOP-constrained environments such as CPUs.
CPU results. Figure 5(a) and 5(b) report the decoding speed under varying prompt lengths on a 4-Core or a 16-Core CPU. We see that with hardware optimization for sparse activation in both Spark FFN and Spark Attention, a speedup that ranges from 1.35x to 1.79x can be achieved on a 16-Core CPU depending on the prompt length. For short prompts (e.g., 256 tokens) and long prompts (e.g. 4096 tokens), the hardware optimization for Spark FFN and Spark Attention provide the most speedup, respectively.
Figure 6: Comparison of training slowdown from using our statistical top-$k$ vs the standard top-$k$ (i.e., jax.lax.approx_max_k [13]) relative to not using any top-$k$.
Figure 1(c) further highlights the efficiency of Spark Transformer in both prefill and decode phases. During the prefill, the prompt is usually chunked into batches since the process is bounded by memory bandwidth. This may reduce the benefit of activation sparsity as different tokens in a chunk may activate different subsets of parameters (in FFN) and attend to different subsets of tokens (in Attention). However, Figure 1(c) shows that Spark Transformer maintains strong performance with a chunk size of 64 tokens, following the default setup in gemma.cpp. A more detailed performance analysis of batching/chunking is provided in Section C.3. In addition, Spark Transformer significantly outperforms Gemma-2 during decoding (with batch size=1).
GPU results. We also evaluate the efficiency gain from sparsity on low-profile GPUs. Figure 5(c) reports the decoding speed under varying prompt lengths on a NVIDIA T4 GPU. Similar to the CPU case, we see Spark Transformer achieves decode speedup ranging from 1.25$\times$ to 1.40$\times$.
@@ 4.4 Training Efficiency
The introduction of a top-$k$ operator is expected to lead to a training slowdown due to the extra computational operators. This slowdown can be prohibitively large if one uses the standard approximate top-$k$ operator provided in JAX, namely jax.lax.approx_max_k [13]. Specifically, this JAX top-$k$ operator is optimized to achieve TPU peak performance and has a controllable recall target, which we vary on the x-axis in Figure 6. It can be observed that the JAX top-$k$ leads to more than 10$\times$ slowdown even when operating on a small recall of 50%. In contrast, the slowdown from statistical top-$k$ is with a very small amount, demonstrating its efficiency. Finally, we do not provide the quality of models trained with JAX top-$k$ since such models take a very long time to train.
@@ 5 Discussion and Related Work
This paper introduces the Spark Transformer architecture to reduce the FLOPs in both the FFN and Attention of Transformer models. Because the FFN and Attention components dominate the computational cost in large Transformers with long contexts, the overall FLOP count for decoding a token is also drastically reduced (see Table C.1), leading to notable computation efficiency benefits on appropriate hardwares. Specifically, the benefit is obtained by selectively activating only part of the model parameters and limiting the attended context for each input. This principle of sparse activation finds a compelling parallel in neuroscience, where studies reveal sparse activity patterns in the brain as a key factor in its remarkable efficiency [5, 7, 44]. While hardware limitations currently hinder the full exploitation of sparse activation, our demonstration of practical wall-time reduction of Spark Transformer on CPU and GPU, together with prior evidence for related techniques [53, 77] including TPU [83], highlight its potential. We hope that this work opens avenues for research into alternative hardware better suited for sparse computations, circumventing the hardware lottery [34] and potentially leading to greater efficiency gains in the future.
@@ 5.1 Towards a mixture of minimum experts?
Mixture of Experts (MoEs) is a prominent class of sparsely activated models which groups the neurons within FFN and activates all neurons in selected groups [73, 48]. This structured grouping facilitates compatibility with modern training accelerators such as GPUs and TPUs, enabling the training of extremely large models. However, reliably training MoEs remains challenging, requiring complex algorithmic design and special hardware support [26]. More crucially, the structured nature of sparsity limits the model’s flexibility and expressiveness, comprising their quality compared to their dense counterparts with an equal parameter count.
To mitigate this, recent work has trended towards architectures with a large number of smaller experts, and activating multiple of them simultaneously [17, 32]. Spark Transformer may be viewed as pushing this trend to the extreme, where each neuron in FFN is an expert itself, with its low-rank predictor vector being the expert router. This gives rise to what may be described as a Mixture of Minimum Experts.
Finally, the discovery of the naturally emerging unstructured activation sparsity has motivated the perspective of naturally emerging experts [85, 22, 15, 65, 80, 89].
@@ 5.2 Reactivating activation sparsity?
Sparse activation is common approach to improve the efficiency of large models and many techniques for a low-cost activation prediction have been developed over the years, such as low-rank factorization [19], quantization [11], product keys [42], hashing [12], etc. With the popularity of modern Transformer models, these techniques become natural choices [36, 84, 53, 77] for reducing their high computation costs. In particular, a lot of the excitement comes from the discovery that the activations in FFNs are naturally sparse [85, 49, 56] and hence efficiency with activation sparsity is obtained without a quality toll.
@@
Our work falls into the category of the latest work in this direction that aims to bring the benefits to the latest generation large language models that do not have natural sparsity. Early attempts [59, 64, 86] seek to bring back sparsity by switching back to ReLU variants, but it usually incurs a quality loss. The quality gap may be largely bridged by more careful tuning, but the activation becomes less sparse (e.g. $25\%$ nonzeros in LLAMA 7B [76]). Top-k has become a more popular choice for obtaining sparsity recently [78] and is able to maintain neutral quality while offering strong sparsity, but only in selected layers [83]. Moreover, such methods require finetuning to bring sparsity and also obtain a predictor. Without finetuning, [45, 51, 88] obtained at most 50% nonzeros under neutral quality. In contrast to these works, our work not only obtains 8% nonzeros in activation of all FFN layers, but also a predictor, all with a single-stage training. We provide a summary of comparison to these methods in Section D.1.
@@
We note that the usefulness of activation sparsity goes beyond efficiency. For example, theoretical studies show its benefits for model generalizability and learnability [60, 6]. Moreover, activated neurons may be associated with semantic concepts, which offers understanding of the working mechanism and enables manipulating the output of Transformer models [16, 55].
@@
Sparse attention broadly refers to the approach of attending to a selected subset of tokens in the context as a means of reducing computation cost [20, 38, 68]. Works on sparse attention include those that use handcrafted attention patterns [14, 9, 2, 21], which feature simplicity, and learned attention patterns [41, 70, 61] which feature better modeling capacity. However, learning attention patterns often involve learning, e.g., a hash table or k-means centers, which significantly complicates modeling. Closely related to our Spark Attention is the top-$k$ attention [31], which obtains data-adaptive attention simply from top-$k$ thresholding. Our work improves upon top-$k$ attention by introducing a low cost predictor which enables an increased computational benefits from sparsity. Finally, KV pruning approaches drop selected tokens permanently as decoding proceeds [87, 52], and cannot achieve as high compression ratio as sparse attention based approaches.
@@ Appendix A Proofs
We first establish the concentration bounds that the empirical mean and standard deviation, i.e., $\bar{x}$ and $s$ are close to the true mean and true standard deviation, i.e., $\mu$ and $\sigma$ of the underlying Gaussian, respectively. Recall from the definition of the chi-squared distribution that $(d-1)\frac{s^{2}}{\sigma^{2}}\sim\chi^{2}(d-1)$. Using the Laurent-Massart bound on the tail probability of the chi-squared distribution [43, Corollary of Lemma 1], we have
@@
for every $t>0$. We set $t=\log\frac{6}{\delta}$. Then, with a probability of at least $31-\delta/$ 3, we have
@@
Equation A.1 provides a concentration bound for $s$. We now proceed to deriving a bound for $\mu$. Towards that, notice that $\frac{\bar{x}-\mu}{\sigma/\sqrt{d}}\sim\mathcal{N}(0,1)$. By using the Mill’s inequality that upper bounds the tail probability of a standard normal distribution (i.e., if $Z\sim\mathcal{N}(0,1)$ and $t>0$, then $\Pr(\left|Z\right|>t)\leq\frac{e^{-t^{2}/2}}{t}$), we have
@@
Therefore, with probability at least $31-\delta/$ 3, we have
@@
Combining Equation A.1 and Equation A.2, with probability at least $31-2\delta/$ 3, we have
@@
$\operatorname{\mathbf{card}}\left(\{i\in[d]~{}|~{}x_{i}>\theta({\bm{x}},k)\}\right)=\sum_{i\in[d]}\bm{1}_{\{x_{i}>\theta({\bm{x}},k)\}}=d\text{$\left(1-\hat{F}_{d}(\theta({\bm{x}},k))\right)\,$.}$
Let $F$ denote the cumulative distribution function (CDF) of $\mathcal{N}(\mu,\sigma^{2})$. By the Dvoretzky-Kiefer-Wolfowitz inequality [25, 58], we have
@@
Applying the union bound on Equation A.3 and Equation A.6, we obtain that the following holds with probability at least $1-\delta$:
@@
To simplify Equation A.7, we consider two cases:
@@
Plugging this into Equation A.7, we obtain
@@
Recall $\operatorname{\mathbf{card}}\left(\{i\in[d]~{}|~{}x_{i}>\theta({\bm{x}},k)\}\right)=d\text{$\left(1-\hat{F}_{d}(\theta({\bm{x}},k))\right)$}.$ We conclude that with probability at least $1-\delta$, we have
$\left|\operatorname{\mathbf{card}}\left(\{i\in[d]~{}|~{}x_{i}>\theta({\bm{x}},k)\}\right)-k\right|\leq 4\sqrt{d\log\frac{6}{\delta}}\left(1+\sqrt{-2\log\min\left\{\frac{k}{d},1-\frac{k}{d}\right\}}\right)\,.$
@@
The Huber statistical top-$k$ in Equation 14 may be written as
@@
where $\theta({\bm{x}},k)$ is defined in Equation 10. This function is the (multivariate) composition of two functions, namely, $\theta=\theta({\bm{x}},k)$ and $\operatorname{Huber}(\operatorname{Soft-Threshold}({\bm{x}},\theta))$. In particular, the former is continuously differentiable (i.e., $C^{1}$) in ${\bm{x}}$, since it is simply a linear combination of sample mean and sample standard deviation both of which are $C^{1}$ functions. To establish the theorem, we only need to show that $\operatorname{Huber}(\operatorname{Soft-Threshold}({\bm{x}},\theta))$ is also a $C^{1}$ function in $({\bm{x}},\theta)$.
@@
$\operatorname{Huber}(\operatorname{Soft-Threshold}(x,\theta))=\begin{cases}\delta x-\delta\theta-\frac{1}{2}\delta,&~{}\text{if}~{}x>\theta+\delta;\\ \frac{1}{2}(x-\theta)^{2},&~{}\text{if}~{}\theta\leq x\leq\theta+\delta;\\ 0,&~{}\text{if}~{}x<\theta.\end{cases}$ (A.14)
@@
$\frac{\partial\operatorname{Huber}(\operatorname{Soft-Threshold}(x,\theta))}{\partial(x,\theta)}=\begin{cases}(\delta,-\delta),&~{}\text{if}~{}x>\theta+\delta;\\ (x-\theta,\theta-x)&~{}\text{if}~{}\theta\leq x\leq\theta+\delta;\\ (0,0),&~{}\text{if}~{}x<\theta,\end{cases}$ (A.15)

### v2

@@ 4.1 Quality
We evaluate Spark Transformer on a suite of benchmarks that are used in the Gemma-2 paper (Gemma Team, 2024), and report the result in Figure˜1(b). We observe that Spark Transformer matches the quality of Gemma-2 while having a drastically reduced FLOP count per token.
The near-quality neutrality of Spark Transformer distinguishes it from related work on enforcing activation sparsity in FFN, which often lead to a quality loss. To demonstrate this, we pretrain variants of Gemma-2 2B with 1) activation function switched to ReLU (Zhang et al., 2024), 2) activation function switched to ReLU2 (Mirzadeh et al., 2023), and 3) Top-k thresholding applied before GELU. Due to high training cost, all models are pretrained for 1/6 of the standard Gemma-2 training iterations. In Figure˜1(a), we report the training losses and FLOPs per token relative to those of the standard Gemma-2. It can be seen that these models either suffer from a large quality loss (i.e., ReLU and Topk), or do not lead to a sufficient FLOPs reduction (i.e., ReLU2). In contrast, our Spark FFN achieves less quality loss with more FLOPs reduction. [^4] Notably, the only difference between Spark FFN and Topk, when both have 8% activated neurons, is in the introduction of the low-cost predictor. Hence, this predictor not only leads to FLOPs reduction but also (surprisingly?) improves model quality. Finally, the combination of Spark FFN with Spark Attention introduces additional FLOPs reduction and quality benefits, enabling an almost neutral quality of Spark Transformer with a large overall FLOPs reduction.
@@ 4.2 Sparsity
To verify the effectiveness of statistical top-$k$, we report the level of sparsity measured in terms of percentage of nonzeros in FFN and the number of nonzeros in Attention. At the beginning of model training, we observe that statistical top-$k$ produces close to 8% nonzeros in FFN (see Figure˜3(a)), which aligns well with our hyper-parameter choice of using $k/d_{\text{ff}}=8\%$ in Spark FFN. This is expected as the model parameters, particularly $\bm{K}$} in Spark FFN, are randomly initialized, hence the entries of the activation maps are drawn from a Gaussian distribution which is in accordance with the assumption of statistical top-$k$. The Gaussian assumption is no longer guaranteed after training, but we empirically observe it to hold approximately (see Section˜C.1) and statistical top-$k$ reliably produce a sparsity level close to 8% until the end of training at $480k$ steps. Sparsity in attention is reported in Figure˜3(b), which show that the number of attended tokens is below our hyper-parameter choice of 256 in Spark Attention throughout training. In particular, the numbers are much smaller because the results are from averaging over all tokens many of which have a context length of less than 256. Finally, we observe comparable levels of sparsity during evaluation (see Section˜C.2).
@@ 4.3 Inference Efficiency
We evaluate the efficiency benefits of Spark Transformer over standard Gemma-2 on both CPUs and GPUs. For CPU evaluation, we use gemma.cpp (Google Gemma.cpp, 2025), the official C++ inference engine optimized for CPUs. For GPU evaluation, we use llama.cpp (Georgi Gerganov, 2025), a widely-used LLM inference engine which supports running a wide selection of LLM models on GPUs. We modify both frameworks to support sparse matrix multiplication operators, which exploit sparsity in both FFN and Attention layers. Our implementation leverages vector SIMD operations (SIMD Wikipedia, 2024) existing in modern CPU, and customized CUDA kernel for GPU. In particular, our implementation reduces not only computational FLOPs but also memory bandwidth requirements, directly accelerating memory-bound scenarios; see Figure˜4 for an illustration and Appendix˜B in Appendix for details. We show that Spark Transformer significantly improves the efficiency of transformer models, even in highly FLOP-constrained environments such as CPUs.
CPU results. Figure˜5(a) and 5(b) report the decoding speed under varying prompt lengths on a 4-Core or a 16-Core CPU. We see that with hardware optimization for sparse activation in both Spark FFN and Spark Attention, a speedup that ranges from 1.35x to 1.79x can be achieved on a 16-Core CPU depending on the prompt length. For short prompts (e.g., 256 tokens) and long prompts (e.g. 4096 tokens), the hardware optimization for Spark FFN and Spark Attention provide the most speedup, respectively.
Figure 6: Comparison of training slowdown from using our statistical top-$k$ vs the standard top-$k$ (i.e., jax.lax.approx_max_k (Chern et al., 2022)) relative to not using any top-$k$.
Figure˜1(c) further highlights the efficiency of Spark Transformer in both prefill and decode phases. During the prefill, the prompt is usually chunked into batches since the process is bounded by memory bandwidth. This may reduce the benefit of activation sparsity as different tokens in a chunk may activate different subsets of parameters (in FFN) and attend to different subsets of tokens (in Attention). However, Figure˜1(c) shows that Spark Transformer maintains strong performance with a chunk size of 64 tokens, following the default setup in gemma.cpp. A more detailed performance analysis of batching/chunking is provided in Section˜C.3. In addition, Spark Transformer significantly outperforms Gemma-2 during decoding (with batch size=1).
[…]
@@ 4.4 Training Efficiency
[…]
@@ 5 Discussion and Related Work
This paper introduces the Spark Transformer architecture to reduce the FLOPs in both the FFN and Attention of Transformer models. Because the FFN and Attention components dominate the computational cost in large Transformers with long contexts, the overall FLOP count for decoding a token is also drastically reduced (see Table˜C.1), leading to notable computation efficiency benefits on appropriate hardwares. Specifically, the benefit is obtained by selectively activating only part of the model parameters and limiting the attended context for each input. This principle of sparse activation finds a compelling parallel in neuroscience, where studies reveal sparse activity patterns in the brain as a key factor in its remarkable efficiency (Attwell and Laughlin, 2001; Barth and Poulet, 2012; Lee et al., 2024a). While hardware limitations currently hinder the full exploitation of sparse activation, our demonstration of practical wall-time reduction of Spark Transformer on CPU and GPU, together with prior evidence for related techniques (Liu et al., 2023; Song et al., 2023) including TPU (Yerram et al., 2024), highlight its potential. We hope that this work opens avenues for research into alternative hardware better suited for sparse computations, circumventing the hardware lottery (Hooker, 2021) and potentially leading to greater efficiency gains in the future.
@@ 5.1 Towards a mixture of minimum experts?
Mixture of Experts (MoEs) is a prominent class of sparsely activated models which groups the neurons within FFN and activates all neurons in selected groups (Shazeer et al., 2017; Lepikhin et al., 2020). This structured grouping facilitates compatibility with modern training accelerators such as GPUs and TPUs, enabling the training of extremely large models. However, reliably training MoEs remains challenging, requiring complex algorithmic design and special hardware support (Fedus et al., 2022). More crucially, the structured nature of sparsity limits the model’s flexibility and expressiveness, comprising their quality compared to their dense counterparts with an equal parameter count.
To mitigate this, recent work has trended towards architectures with a large number of smaller experts, and activating multiple of them simultaneously (Dai et al., 2024; He, 2024). Spark Transformer may be viewed as pushing this trend to the extreme, where each neuron in FFN is an expert itself, with its low-rank predictor vector being the expert router. This gives rise to what may be described as a Mixture of Minimum Experts.
Finally, the discovery of the naturally emerging unstructured activation sparsity has motivated the alternative perspective of designing experts by grouping neurons from their natural sparsity (Zhang et al., 2022; Dong et al., 2023; Csordás et al., 2023; Qiu et al., 2023; Szatkowski et al., 2024; Zheng et al., 2024).
@@ 5.2 Reactivating activation sparsity?
Sparse activation is common approach to improve the efficiency of large models and many techniques for a low-cost activation prediction have been developed over the years, such as low-rank factorization (Davis and Arel, 2013), quantization (Cao et al., 2019), product keys (Lample et al., 2019), hashing (Chen et al., 2020), etc. With the popularity of modern Transformer models, these techniques become natural choices (Jaszczur et al., 2021; Zeng et al., 2023; Liu et al., 2023; Song et al., 2023) for reducing their high computation costs. In particular, a lot of the excitement comes from the discovery that the activations in FFNs are naturally sparse (Zhang et al., 2022; Li et al., 2022; Luo et al., 2024b) and hence efficiency with activation sparsity is obtained without a quality toll.
@@
Our work falls into the category of the latest work in this direction that aims to bring the benefits to the latest generation large language models that do not have natural sparsity. Early attempts (Mirzadeh et al., 2023; Peng et al., 2023; Zhang et al., 2024) seek to bring back sparsity by switching back to ReLU variants, but it usually incurs a quality loss. The quality gap may be largely bridged by more careful tuning, but the activation becomes less sparse (e.g. $25\%$ nonzeros in LLAMA 7B (Song et al., 2024a)). Top-k has become a more popular choice for obtaining sparsity recently (Song et al., 2024b) and is able to maintain neutral quality while offering strong sparsity, but only in selected layers (Yerram et al., 2024). Moreover, such methods require finetuning to bring sparsity and also obtain a predictor. Without finetuning, Lee et al. (2024b); Liu et al. (2024a); Zhang et al. (2025) obtained at most 50% nonzeros under neutral quality. In contrast to these works, our work not only obtains 8% nonzeros in activation of all FFN layers, but also a predictor, all with a single-stage training. We provide a summary of comparison to these methods in Section˜D.1.
@@
We note that the usefulness of activation sparsity goes beyond efficiency. For example, theoretical studies show its benefits for model generalizability and learnability (Muthukumar and Sulam, 2023; Awasthi et al., 2024). Moreover, activated neurons may be associated with semantic concepts, which offers understanding of the working mechanism and enables manipulating the output of Transformer models (Cuadros et al., 2022; Luo et al., 2024a).
@@
Sparse attention broadly refers to the approach of attending to a selected subset of tokens in the context as a means of reducing computation cost (Deng et al., 2024; Jiang et al., 2024; Ren et al., 2023). Works on sparse attention include those that use handcrafted attention patterns (Child et al., 2019; Beltagy et al., 2020; Ainslie et al., 2023; Ding et al., 2023), which feature simplicity, and learned attention patterns (Kitaev et al., 2020; Roy et al., 2021; Pagliardini et al., 2023) which feature better modeling capacity. However, learning attention patterns often involve learning, e.g., a hash table or k-means centers, which significantly complicates modeling. Closely related to our Spark Attention is the top-$k$ attention (Gupta et al., 2021), which obtains data-adaptive attention simply from top-$k$ thresholding. Our work improves upon top-$k$ attention by introducing a low cost predictor which enables an increased computational benefits from sparsity. Finally, KV pruning approaches drop selected tokens permanently as decoding proceeds (Zhang et al., 2023; Liu et al., 2024b), and cannot achieve as high compression ratio as sparse attention based approaches.
5.3 Synergizing for greater inference efficiency?
This work also opens promising avenues for future research on efficient inference of large language models, particularly in combining Spark Transformer’s architectural efficiency with other leading optimization techniques. We briefly discuss two key synergies here.
Synergy with Speculative Decoding. Spark Transformer is highly complementary to speculative decoding. As a target model, its faster inference directly accelerates the primary verification bottleneck. As a draft model, its near-quality neutrality and high speed make it an ideal candidate for generating high-quality drafts, potentially leading to higher token acceptance rates and greater overall speedups. A full discussion is provided in Section˜D.3.
Synergy with Quantization. We also hypothesize a strong synergy with quantization. The benefits are expected to be multiplicative, as Spark Transformer reduces the number of operations while quantization reduces their cost. More importantly, unlike standard pruning which preserves high-magnitude outliers, our statistical top-$k$ operator uses soft-thresholding. This shrinks the dynamic range of activation distribution, which may reduce sensitivity to activation quantization. A detailed analysis of this mechanism is available in Appendix section˜D.4.
@@ Appendix A Proofs
We first establish the concentration bounds that the empirical mean and standard deviation, i.e., $\bar{x}$ and $s$ are close to the true mean and true standard deviation, i.e., $\mu$ and $\sigma$ of the underlying Gaussian, respectively. Recall from the definition of the chi-squared distribution that $(d-1)\frac{s^{2}}{\sigma^{2}}\sim\chi^{2}(d-1)$. Using the Laurent-Massart bound on the tail probability of the chi-squared distribution (Laurent and Massart, 2000, Corollary of Lemma 1), we have
@@
for every $t>0$. We set $t=\log\frac{6}{\delta}$. Then, with a probability of at least $1-\delta/3$, we have
@@
Equation˜A.1 provides a concentration bound for $s$. We now proceed to deriving a bound for $\mu$. Towards that, notice that $\frac{\bar{x}-\mu}{\sigma/\sqrt{d}}\sim\mathcal{N}(0,1)$. By using the Mill’s inequality that upper bounds the tail probability of a standard normal distribution (i.e., if $Z\sim\mathcal{N}(0,1)$ and $t>0$, then $\Pr(\left|Z\right|>t)\leq\frac{e^{-t^{2}/2}}{t}$), we have
@@
Therefore, with probability at least $1-\delta/3$, we have
@@
Combining Equation˜A.1 and Equation˜A.2, with probability at least $1-2\delta/3$, we have
@@
$\operatorname{\mathbf{card}}\left(\{i\in[d]\penalty 10000\ |\penalty 10000\ x_{i}>\theta({\bm{x}},k)\}\right)=\sum_{i\in[d]}\bm{1}_{\{x_{i}>\theta({\bm{x}},k)\}}=d\text{$\left(1-\hat{F}_{d}(\theta({\bm{x}},k))\right)\,$.}$
Let $F$ denote the cumulative distribution function (CDF) of $\mathcal{N}(\mu,\sigma^{2})$. By the Dvoretzky-Kiefer-Wolfowitz inequality (Dvoretzky et al., 1956; Massart, 1990), we have
@@
Applying the union bound on Equation˜A.3 and Equation˜A.6, we obtain that the following holds with probability at least $1-\delta$:
@@
To simplify Equation˜A.7, we consider two cases:
@@
Plugging this into Equation˜A.7, we obtain
@@
Recall $\operatorname{\mathbf{card}}\left(\{i\in[d]\penalty 10000\ |\penalty 10000\ x_{i}>\theta({\bm{x}},k)\}\right)=d\text{$\left(1-\hat{F}_{d}(\theta({\bm{x}},k))\right)$}.$ We conclude that with probability at least $1-\delta$, we have
$\left|\operatorname{\mathbf{card}}\left(\{i\in[d]\penalty 10000\ |\penalty 10000\ x_{i}>\theta({\bm{x}},k)\}\right)-k\right|\leq 4\sqrt{d\log\frac{6}{\delta}}\left(1+\sqrt{-2\log\min\left\{\frac{k}{d},1-\frac{k}{d}\right\}}\right)\,.$
@@
The Huber statistical top-$k$ in Equation˜14 may be written as
@@
where $\theta({\bm{x}},k)$ is defined in Equation˜10. This function is the (multivariate) composition of two functions, namely, $\theta=\theta({\bm{x}},k)$ and $\operatorname{Huber}(\operatorname{Soft-Threshold}({\bm{x}},\theta))$. In particular, the former is continuously differentiable (i.e., $C^{1}$) in $\bm{x}$}, since it is simply a linear combination of sample mean and sample standard deviation both of which are $C^{1}$ functions. To establish the theorem, we only need to show that $\operatorname{Huber}(\operatorname{Soft-Threshold}({\bm{x}},\theta))$ is also a $C^{1}$ function in $({\bm{x}},\theta)$.
@@
$\operatorname{Huber}(\operatorname{Soft-Threshold}(x,\theta))=\begin{cases}\delta x-\delta\theta-\frac{1}{2}\delta,&\penalty 10000\ \text{if}\penalty 10000\ x>\theta+\delta;\\ \frac{1}{2}(x-\theta)^{2},&\penalty 10000\ \text{if}\penalty 10000\ \theta\leq x\leq\theta+\delta;\\ 0,&\penalty 10000\ \text{if}\penalty 10000\ x<\theta.\end{cases}$ (A.14)
@@
$\frac{\partial\operatorname{Huber}(\operatorname{Soft-Threshold}(x,\theta))}{\partial(x,\theta)}=\begin{cases}(\delta,-\delta),&\penalty 10000\ \text{if}\penalty 10000\ x>\theta+\delta;\\ (x-\theta,\theta-x)&\penalty 10000\ \text{if}\penalty 10000\ \theta\leq x\leq\theta+\delta;\\ (0,0),&\penalty 10000\ \text{if}\penalty 10000\ x<\theta,\end{cases}$ (A.15)